OpenStudy (studygurl14):

Evaluate? help @amistre64 @ganeshie8 @SolomonZelman @Compassionate

OpenStudy (studygurl14):

\(\large\cos(\arctan\frac{3}{4})\)

OpenStudy (studygurl14):

Is it 4/5?

OpenStudy (studygurl14):

Evaluate \(\large\cos(2\pi+\arctan\frac{4}{3})\)

OpenStudy (solomonzelman):

the first one, yes is 4/5.

OpenStudy (solomonzelman):

the second one too. 4/5

OpenStudy (studygurl14):

The second one too? Can you show me how?

OpenStudy (studygurl14):

@SolomonZelman

OpenStudy (solomonzelman):

Ohh, I did, arctan(3/4) instead of 4/3, lol (for the second one)

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ \cos\left\{ 2\pi+\arctan( \frac{4}{3}) \right\}=\cos(2\pi ) \cos\left\{ \arctan( \frac{4}{3}) \right\}-\sin(2\pi ) \sin\left\{ \arctan( \frac{4}{3}) \right\} }\)

OpenStudy (solomonzelman):

Using the cos(a+b) rule. Tell me what you get after simplifying this expression?

OpenStudy (studygurl14):

umm....?

OpenStudy (solomonzelman):

Well, you know that \(\normalsize\color{black}{ \cos(2\pi)=-1 }\) and \(\normalsize\color{black}{ \sin(2\pi)=0 }\)

OpenStudy (solomonzelman):

If you need help, tag me back into the question.

OpenStudy (studygurl14):

Simplified to this? \(\large -\cos({\arctan\frac{4}{3}})\)

OpenStudy (solomonzelman):

yes;)

OpenStudy (studygurl14):

omg, thanks so much. I got -0.6

OpenStudy (solomonzelman):

yes:) that's correct!

OpenStudy (studygurl14):

wait..could it be 0.6? cause -0.6 isn't an answer choice.

OpenStudy (solomonzelman):

It is -0.6, because the `cos(2π)` by which you multiplied, is -1. So it should be -0.6, not positive. I am failry sure.

OpenStudy (solomonzelman):

that in a box, is cos 2 pi.

OpenStudy (studygurl14):

maybe the answer choice forgot the minus sign...

OpenStudy (solomonzelman):

well, that is what you should talk to your teach about. Or maybe some info is missling. Like it has to be in first quadrant or something like this?

OpenStudy (studygurl14):

idk...questions at my school are really poorly written

OpenStudy (solomonzelman):

but this is all you had? To evaluate, cos( 2pi + arctan(4/3) ) /

OpenStudy (solomonzelman):

(the slash shouldn't be there)

OpenStudy (studygurl14):

yes, it is

OpenStudy (studygurl14):

"evaluate" and then that equation